# Bridging the Rossby number gap in rapidly rotating thermal convection

### v1

@@
Bridging the Rossby number gap in rapidly rotating thermal convection
Adrian van Kan,1† Keith Julien,2111Deceased on 14th of April 2024.   Benjamin Miquel,3 Edgar Knobloch1
1 Department of Physics, University of California at Berkeley, Berkeley, CA 94720, USA
2 Department of Applied Mathematics, University of Colorado, Boulder, CO 80309, USA
3 CNRS, École Centrale de Lyon, INSA Lyon, Université Claude Bernard Lyon 1,
Laboratoire de Mécanique des Fluides et d’Acoustique, UMR5509, F-69134 Écully, France
To whom correspondence should be addressed; E-mail: avankan@berkeley.edu
@@ 3 CNRS, École Centrale de Lyon, INSA Lyon, Université Claude Bernard Lyon 1,
Geophysical and astrophysical fluid flows are typically buoyantly driven and are strongly constrained by planetary rotation at large scales. Rapidly rotating Rayleigh-Bénard convection (RRRBC) provides a paradigm for direct numerical simulations (DNS) and laboratory studies of such flows, but the accessible parameter space remains restricted to moderately fast rotation (Ekman numbers ${\rm Ek}\gtrsim 10^{-8}$), while realistic ${\rm Ek}$ for astro-/geophysical applications are significantly smaller. Reduced equations of motion, the non-hydrostatic quasi-geostrophic equations describing the leading-order behavior in the limit of rapid rotation (${\rm Ek}\to 0$) cannot capture finite rotation effects, leaving the physically most relevant part of parameter space with small but finite ${\rm Ek}$ currently inaccessible. Here, we introduce the rescaled incompressible Navier-Stokes equations (RiNSE) – a reformulation of the Navier-Stokes-Boussinesq equations informed by the scalings valid for ${\rm Ek}\to 0$. We provide the first full DNS of RRRBC at unprecedented rotation strengths down to $\rm Ek=10^{-15}$ and below, and show that the RiNSE converge to the asymptotically reduced equations.
Introduction
@@
The universe abounds with examples of turbulent flows that are driven by buoyancy and constrained by rotation. These highly complex flows shape our environment across scales: in protoplanetary disks, they mediate the formation of planetesimals  (?); in planetary atmospheres, including on Earth, they carry heat, momentum and moisture, crucially impacting the climate  (?, ?, ?); in the Earth’s oceans  (?, ?), they transport heat, carbon dioxide, and nutrients, thereby crucially influencing biotopes  (?); in planetary and stellar cores, they are responsible for the generation of large-scale magnetic fields via the dynamo instability  (?, ?, ?); in the subsurface oceans of the gaseous giants’ moons, they shape the icy crusts of Europa, Enceladus, Titan, etc.  (?, ?, ?, ?); in stars like the Sun  (?, ?) they are responsible for large-scale magnetic fields as well as rich dynamics such as the 22-year cycle in solar activity.
@@
Such geo- and astrophysical flows, in spite of their differences, typically share two important properties. Owing to their turbulent nature, their dynamics result from a broad range of interacting spatial and temporal scales. Accordingly, advection by the flow dominates over viscous forces when the Reynolds number is large, ${\rm Re}_{H}=UH/\nu\gg 1$, based here on the fluid layer depth $H$, a typical velocity of the flow $U$, and kinematic viscosity $\nu$. Simultaneously, advection by the flow occurs on time scales much longer than the planetary rotation period, as measured by small Rossby numbers ${\rm Ro}_{H}=U/(2\Omega H)\ll 1$, where $\Omega$ is the planetary rotation rate.    Also of particular importance is the ratio of the viscous and Coriolis forces providing an a priori external parameter referred to as the Ekman number ${\rm Ek}={\rm Ro}_{H}/{\rm Re}_{H}=\nu/(2\Omega H^{2})$, as well as the Prandtl number $\sigma=\nu/\kappa$ characterizing the fluid under consideration, where $\kappa$ is the thermal diffusivity. The strength of the thermal forcing is controlled by the Rayleigh number ${\rm Ra}$, which is proportional to the temperature drop $\Delta_{T}$ across the fluid layer (see Methods and Materials section).
@@
Table 1: Nondimensional parameter estimates for planetary  (?), satellite  (?) and stellar interiors  (?, ?). Estimates of the Rossby number are derived from the relation $\rm Ro_{H}=Re_{H}Ek$.
@@
The dynamics of the fluid flows found in the celestial bodies listed in Table 1 are further rendered complex by additional ingredients such as spherical geometry, multiple contributors to the density, compressibility, and the presence of magnetic fields. Absent such complexities, the quintessential paradigm for investigating rotationally influenced buoyant flows is provided by rapidly rotating Rayleigh-Bénard convection (RRRBC). A large number of studies have been published on this model system, which is very well suited for detailed experimental, numerical and theoretical studies, see for example  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and the recent review  (?). In RRRBC, gravity is often taken to be antiparallel to the rotation on the assumption that the Froude number ${\rm Fr}_{\Omega}\equiv\Omega^{2}H/g$ is small. The resulting configuration is translation-invariant and believed to be relevant to the polar regions of planets, moons and stars as well as to their interiors, although the case of misaligned gravity and rotation axis has also been studied  (?, ?, ?, ?, ?).
@@
A major obstacle in the study of rapidly rotating convection, even in the highly idealized setting of RRRBC, is the difficulty, to date, of attaining the extremely low Ekman numbers, typically $10^{-12}$ or smaller (cf. Table 1), that characterize geo- and astrophysical flows. By contrast, current state-of-the-art DNS of the full Navier-Stokes-Boussinesq equations that describe RRRBC have been restricted to ${\rm Ek}\gtrsim 10^{-8}$  (?, ?, ?, ?), despite enormous computational efforts. Inclusion of magnetic effects in the simulations to capture the capability for dynamo action, as well as spherical shell geometry, further limits the reported range of nearly all studies to ${\rm Ek}\gtrsim 10^{-7}$, ${\rm Re}_{H}\lesssim 10^{3}$  (?, ?, ?, ?, ?), although $\rm Ek=10^{-8}$ has been achieved in this context (?) but at great computational expense; simulations down to ${\rm Ek}=10^{-10}$ are available for linearized dynamics only  (?). In state-of-the-art laboratory experiments, the accessible parameter space is even more restricted, to ${\rm Ek}\gtrsim 10^{-7}$ (?, ?, ?, ?, ?). These accessible parameter values in numerical or laboratory experiments are orders of magnitude larger than the estimates for realistic geo-/astrophysical applications. On the other hand, existing reduced equations  (?), known as the non-hydrostatic quasi-geostrophic (NHQG) equations, describe the dynamics in the limit ${\rm Ek}\to 0$, but their applicability is by construction confined to asymptotically small values of ${\rm Ek}$. This leaves an unexplored gap corresponding to small but finite Ekman numbers, precisely the regime that is most relevant for geophysical and astrophysical applications.
@@
Results

### v2

@@
Abstract
@@
Geophysical and astrophysical fluid flows are typically driven by buoyancy and strongly constrained at large scales by planetary rotation. Rapidly rotating Rayleigh-Bénard convection (RRRBC) provides a paradigm for experiments and direct numerical simulations (DNS) of such flows, but the accessible parameter space remains restricted to moderately fast rotation rates (Ekman numbers ${\rm Ek}\gtrsim 10^{-8}$), while realistic ${\rm Ek}$ for astro-/geophysical applications are orders of magnitude smaller. On the other hand, previously derived reduced equations of motion describing the leading-order behaviour in the limit of very rapid rotation (${\rm Ek}\to 0$) cannot capture finite rotation effects, and the physically most relevant part of parameter space with small but finite ${\rm Ek}$ has remained elusive. Here, we employ the rescaled incompressible Navier-Stokes equations (RiNSE) – a reformulation of the Navier-Stokes-Boussinesq equations informed by the scalings valid for ${\rm Ek}\to 0$, recently introduced by Julien et al. (2024b) – to provide full DNS of RRRBC at unprecedented rotation strengths down to $\rm Ek=10^{-15}$ and below, revealing the disappearance of cyclone-anticyclone asymmetry at previously unattainable Ekman numbers (${\rm Ek}\approx 10^{-9}$). We also identify an overshoot in the heat transport as ${\rm Ek}$ is varied at fixed $\rm\widetilde{Ra}=Ra{\rm Ek}^{4/3}$, associated with dissipation due to ageostrophic motions in the boundary layers. The simulations validate theoretical predictions based on thermal boundary layer theory for RRRBC and show that the solutions of RiNSE agree with the reduced equations at very small ${\rm Ek}$. These results represent a first foray into the vast, largely unexplored parameter space of very rapidly rotating convection rendered accessible by RiNSE.
keywords:
to be added in proof
MSC Codes 76U05, 76F35
1 Introduction
@@
The universe abounds with examples of turbulent flows that are driven by buoyancy and constrained by rotation. These highly complex flows shape our environment across scales: in protoplanetary disks, they mediate the formation of planetesimals (Lesur & Ogilvie, 2010); in planetary atmospheres, including on Earth, they carry heat, momentum and moisture, crucially impacting the climate (Emanuel et al., 1994; Siegelman et al., 2022; Dauxois et al., 2021); in the Earth’s oceans (Marshall & Schott, 1999; Cheon & Gordon, 2019), they transport heat, carbon dioxide, and nutrients, thereby crucially influencing biotopes (Severin et al., 2014); in planetary and stellar cores, they are responsible for the generation of large-scale magnetic fields via the dynamo instability (King et al., 2010; Jones, 2011; Schaeffer et al., 2017); in the subsurface oceans of the gaseous giants’ moons, they shape the icy crusts of Europa, Enceladus, Titan, etc. (Pappalardo et al., 1998; Mitri & Showman, 2008; Nimmo & Pappalardo, 2016; Soderlund et al., 2024); in stars like the Sun (Miesch, 2000; Fan, 2021) they are responsible for large-scale magnetic fields as well as rich dynamics such as the 22-year cycle in solar activity.
@@
Such geo- and astrophysical flows, in spite of their differences, typically share two important properties. Owing to their turbulent nature, their dynamics result from a broad range of interacting spatial and temporal scales. Accordingly, advection by the flow dominates over viscous forces when the Reynolds number is large, ${\rm Re}_{H}=UH/\nu\gg 1$, based here on the fluid layer depth $H$, a typical velocity of the flow $U$, and kinematic viscosity $\nu$. Simultaneously, advection by the flow occurs on time scales much longer than the planetary rotation period, as measured by small Rossby numbers ${\rm Ro}_{H}=U/(2\Omega H)\ll 1$, where $\Omega$ is the planetary rotation rate.
Also of particular importance is the ratio of the viscous and Coriolis forces providing an a priori external parameter referred to as the Ekman number ${\rm Ek}={\rm Ro}_{H}/{\rm Re}_{H}=\nu/(2\Omega H^{2})$, as well as the Prandtl number $\sigma=\nu/\kappa$ characterizing the fluid under consideration, where $\kappa$ is the thermal diffusivity. The strength of the thermal forcing is controlled by the Rayleigh number ${\rm Ra}$, which is proportional to the temperature drop $\Delta_{T}$ across the fluid layer.
@@
Table 1: Nondimensional parameter estimates for planetary (Schubert & Soderlund, 2011), satellite (Soderlund, 2019) and stellar interiors (Miesch, 2000; Garaud & Garaud, 2008). Estimates of the Rossby number are derived from the relation $\rm Ro_{H}=Re_{H}Ek$.
@@
The dynamics of the fluid flows found in the celestial bodies listed in Table 1 are further rendered complex by additional ingredients such as spherical geometry, multiple contributors to the density, compressibility, and the presence of magnetic fields. Absent such complexities, the quintessential paradigm for investigating rotationally influenced buoyant flows is provided by rapidly rotating Rayleigh-Bénard convection (RRRBC). A large number of studies have been published on this model system, which is very well suited for detailed experimental, numerical and theoretical studies, see for example (Chandrasekhar, 1953; Nakagawa & Frenzen, 1955; Veronis, 1959; Rossby, 1969; Boubnov & Golitsyn, 1986; Zhong et al., 1991; Julien et al., 1996; Knobloch, 1998; Hart et al., 2002; Vorobieff & Ecke, 2002; Boubnov & Golitsyn, 2012; King et al., 2012; Stevens et al., 2013) and the recent review by Ecke & Shishkina (2023). In RRRBC, gravity is often taken to be antiparallel to the rotation on the assumption that the Froude number ${\rm Fr}_{\Omega}\equiv\Omega^{2}H/g$ is small. The resulting configuration is translation-invariant and believed to be relevant to the polar regions of planets, moons and stars as well as to their interiors, although the case of misaligned gravity and rotation axis has also been studied (Hathaway & Somerville, 1983; Julien & Knobloch, 1998; Julien et al., 2006; Miquel et al., 2018; Novi et al., 2019; Currie et al., 2020; Cai, 2020; Julien et al., 2024a).
@@
A major obstacle in the study of rapidly rotating convection, even in the highly idealized setting of RRRBC, is the difficulty, to date, of attaining the extremely low Ekman numbers, typically $10^{-12}$ or smaller (cf. Table 1), that characterize geo- and astrophysical flows. By contrast, current state-of-the-art DNS of the full Navier-Stokes-Boussinesq equations that describe RRRBC have been restricted to ${\rm Ek}\gtrsim 10^{-8}$ (Kunnen et al., 2016; Guervilly et al., 2019; Song et al., 2024b, a), despite enormous computational efforts. Inclusion of magnetic effects in the simulations to capture the capability for dynamo action, as well as spherical shell geometry, further limits the reported range of nearly all studies to ${\rm Ek}\gtrsim 10^{-7}$, ${\rm Re}_{H}\lesssim 10^{3}$ (Schaeffer et al., 2017; Cooper et al., 2020; Mason et al., 2022; Kolhey et al., 2022; Majumder et al., 2024; Gastine & Aurnou, 2023), although $\rm Ek=10^{-8}$ has been achieved in this context (Aubert et al., 2017) but at great computational expense; simulations down to ${\rm Ek}=10^{-10}$ are available for linearized dynamics only (He et al., 2022). In state-of-the-art laboratory experiments, the accessible parameter space is even more restricted, to ${\rm Ek}\gtrsim 10^{-7}$(Shew & Lathrop, 2005; Aurnou et al., 2015; Rajaei et al., 2017; Lu et al., 2021; Hawkins et al., 2023; Pothérat & Horn, 2024). These accessible parameter values in numerical or laboratory experiments are orders of magnitude larger than the estimates for realistic geo- and astrophysical applications. On the other hand, existing reduced equations (Sprague et al., 2006), known as the non-hydrostatic quasi-geostrophic (NHQG) equations, describe the dynamics in the limit ${\rm Ek}\to 0$, but their applicability is, by construction, confined to asymptotically small values of ${\rm Ek}$. This leaves an unexplored gap corresponding to small but finite Ekman numbers, precisely the regime that is most relevant for geo- and astrophysical applications.
@@
In the following section we describe the RiNSE and sketch the numerical method used to solve them, followed in section 3 by a description of the results obtained, focusing both on bulk transport properties and the structure of the boundary layers at the top and bottom. The paper concludes with a discussion of the significance of the RiNSE reformulation and prospects for future work. An Appendix provides additional details on the flow statistics and structure.
2 Methodology
In this section, we succinctly present the rescaled Navier-Stokes equations (RiNSE), an equivalent reformulation of the full Navier-Stokes-Boussinesq equations, informed by the scaling behavior valid in the limit of vanishing ${\rm Ek}$ (Julien & Knobloch, 1998; Julien et al., 2012b). The RiNSE formulation is described in detail in a recent paper (Julien et al., 2024b), where it was first introduced and validated against existing results in the literature, and where the full details of the numerical implementation are given.